OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 7x – 5y = 28? Choose all answers that are correct. (−6, −14) (−1, −7) (4, 10) (7, 9)

OpenStudy (texaschic101):

well lets see.. 7x - 5y = 28 1) (-6,-14)...x = -6 and y = -14 now we sub 7(-6) - 5(-14) = 28 -42 + 70 = 28 28 = 28 (correct) --- this one is a solution 2) (-1, -7)...x = -1 and y = -7 now we sub 7(-1) - 5(-7) = 28 -7 + 35 = 28 28 = 28 (correct) -- this one works as well do you see how I am doing this ? Do you understand How to do it now ? I will leave the last two for you to figure out. Just reember, that at the end, if they do not equal, it is not a solution. If you have any questions, just ask